# Sequence to Sequence Reward Modeling: Improving RLHF by Language Feedback

### v1

@@ Introduction
Large Language Models (LLMs) (Achiam et al. 2023; Touvron et al. 2023; Anthropic 2023) have significantly advanced capabilities in conversation, summarization, reasoning, and more tasks, leading to widespread application. Alignment techniques (Ji et al. 2023; Anwar et al. 2024; Bai et al. 2022), particularly Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Christiano et al. 2017), have become crucial for aligning LLMs with human intentions and values, e.g., reducing toxicity (Bai et al. 2022; Ganguli et al. 2022). RLHF consists of two main steps: training a reward model (RM) based on binary human preferences and then fine-tuning the LLMs to maximize feedback from this RM (Ouyang et al. 2022; Schulman et al. 2017; Scheurer et al. 2023). Current reward modeling implementations typically use the sequence-to-scalar (seq2scalar) approach, where a sequence, combined with a prompt and response, is scored as a single scalar value (Huang et al. 2024).
Figure 1: The traditional sequence-to-scalar reward model provides only coarse-grained scalar feedback on the LLM’s response, making it prone to exploiting unexpected generalization for high rewards during the RL fine-tuning phase (Skalse and Abate 2023), e.g., falling into a refusal-to-answer paradigm (Bai et al. 2022; Dai et al. 2023). Without additional annotations, we propose a novel sequence-to-sequence reward modeling method that offers richer language feedback, improving RLHF performance.
@@
However, Due to high divergence, noise, and sparsity in crowd-annotated preference datasets (Wu et al. 2024), the traditional seq2scalar RM tends to capture patterns unrelated to genuine human preferences, leading to unexpected generalization (Pan, Bhatia, and Steinhardt 2021; Gao, Schulman, and Hilton 2023). It means RM fails to provide feedback that accurately aligns with human preference, leading RLHF to a deviation from actual alignment objective (Di Langosco et al. 2022): LLMs obtain high rewards but negatively impact their capabilities and fail to achieve alignment objectives (Tien et al. 2022; Knox et al. 2023; Lambert and Calandra 2023). For example, when the alignment objective is to reduce toxic responses, LLMs may learn the refusal-to-response paradigm instead of providing safe and helpful responses (Dai et al. 2023; Bai et al. 2022). Existing methods require additional annotations to provide more comprehensive feedback, including regularization terms (Moskovitz et al. 2023; Zhu, Jordan, and Jiao 2024), models ensembling (Coste et al. 2023), parameters combinations (Ramé et al. 2024) and additional model (Dai et al. 2023), raising concerns about their robustness and scalability.
@@
  • Our further experiments demonstrate that the seq2seq RM not only mitigates unexpected behavior of RLHF but also enhances its alignment performance. Specifically, it achieves an average improvement of 76.9% in 3 NLP tasks, across models with 2B and 7B parameters.
@@
Given an LLM parameterized with $\theta$, we model its generation as a token-level Markov Decision Process (MDP), i.e., the tuple $\mathcal{T}=(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P})$. Here, let $t\in\mathcal{V}$ denote a token, where $\mathcal{V}$ represents the vocabulary set, and $L$ denotes the maximum number of tokens. The state space $\mathcal{S}$ can be defined as $\{(t_{1},t_{2},\ldots,t_{L})\mid t_{i}\in\mathcal{V}\}$. The action space $\mathcal{A}=\mathcal{V}$ consists of all possible tokens in vocabulary. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns a reward value for generating a specific token given the current sequence of tokens. The transition probability function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the deterministic state transition model, concatenating the generated token to the existing sequence to form the next state. i.e., $s_{t+1}=(s_{t},a_{t})$.
@@
$\mathcal{L}_{\bm{\theta}}^{\text{RL}}=$ $-\mathbb{E}_{a_{t}\sim\pi_{\bm{\theta}}(\cdot|s_{t})}$
$[\sum_{t=0}^{T}(\mathcal{R}(s_{t},a_{t})+\beta\pi_{ref}(a_{t}|s_{t}))]$
where $\pi_{ref}$ is the reference distribution from the initial LLM, combined with coefficient $\beta$ to compute the KL divergence, preventing the actor LLM from deviating too much from the $\pi_{ref}$.
@@ Methods
The seq2scalar RM maps concatenated prompt and response pairs, denoted as $\bm{x},\bm{y^{l}}$ and $\bm{x},\bm{y^{w}}$, to distinct points in a scalar space. This makes it difficult for the token-level differences between responses to be distinguished in the scalar output. To enhance RM’s ability to discern fine-grained differences in responses, we propose augmenting this mapping from a scalar to a language space. Specifically, we extend the learning process from binary MLE in Equation 1 to SFT, a sequence MLE approach. This extension transforms the learning objective from scalar classification to sequence generation, achieving a higher distinction of responses, as shown in Equation Reward Modeling as Sequence MLE. Similar approaches have also been applied by Ji et al. (2024a) to make LLM aligners.
@@
Since the language feedback provided by the seq2seq reward model (RM) cannot be directly used for reinforcement learning (RL) optimization, we design a novel reward extraction mechanism that returns positive and negative scores as token-level feedback. Formally, let $\bm{x}$ denote the prompt used in RLHF, and $\bm{y}$ is the response generated by the actor LLM. We set $j$ as the first index where $r^{\text{seq}}(\bm{x},\bm{y})$ and $\bm{y}$ are not equal, with the subscript indicating the element at the corresponding position. The notation $:j$ indicates the subsequence up to that index. Defining the inference function of the seq2seq RM as $r^{\text{seq}}(\cdot)$, our goal is to leverage $r^{\text{seq}}(\cdot)$ to obtain the token-level score $\mathcal{R}(\bm{x},\bm{y_{:i}})$ for all subsequence $\bm{y_{:i}}$ in $\bm{y}$. To achieve this, we first concatenate $\bm{x}$ and $\bm{y}$ and then input $\bm{y}$ token by token into the seq2seq RM. We positively reward tokens that are consistent with the seq2seq RM at the beginning and negatively reward those that become inconsistent thereafter. Figure 2 provides an intuitive example of this process.
@@
$\mathcal{R}(\bm{x},\bm{y_{:i-1}})=\begin{cases}1&\text{if }r^{\text{seq}}(\bm{x},\bm{y},\bm{y_{:i-1})}=\bm{y_{i}}\text{ and }i<j\\ -1&\text{otherwise}.\end{cases}$
@@
We conducted experiments on the popular open-source LLMs Gemma-2B (Team et al. 2024) and Llama2-7B (Touvron et al. 2023). The selected tasks included single-turn safe dialogue, dialogue summarization, and text summarization. For single-turn safe dialogue, we chose the PKU-SafeRLHF dataset (Ji et al. 2024b). To verify the performance of our method on human preference and synthetic preference settings, we used the Dialogusum (Chen et al. 2021) for the dialogue summarization task and the TL;DR (Stiennon et al. 2020; Völske et al. 2017) for text summarization. For the former, we employed GPT-4 for preference annotation, while the latter already contains human-annotated preferences.
@@
Our evaluation primarily uses GPT-4 evaluation and phenomenon analysis for specific tasks. In single-turn safety dialogue tasks, we analyze whether LLMs fall into an overly conservative refusal-to-response paradigm, leading to a significant drop in utility. Therefore, we use the helpfulness reward model scores provided by Dai et al. (2023) an MT Bench score (Zheng et al. 2024) as the metric of utility. In text summarization and dialogue summarization tasks, we refer to (Stiennon et al. 2020) and use accuracy, conciseness, and information consistency as evaluation criteria. To avoid randomness, we ran the above evaluation over 3 times.
@@
Our experiment includes comparisons between popular baselines and our proposed variants. We define their abbreviations as follows. Baselines. Init-SFT: Pre-trained model finetuned for the corresponding task, also used as the initial model for other alignment methods. PPO: Vanilla PPO trained based on seq2scalar RM (Ouyang et al. 2022; Schulman et al. 2017). DPO: Direct Preference Optimization (Rafailov et al. 2024b). RFT: Rejection sampling fine-tuning (Touvron et al. 2023) based on seq2scalar RM selection over 8 responses generated by the Init-SFT. Variants. PPO-T: PPO trained based on seq2seq RM, with token-level reward. RFT-S: RFT based on seq2seq RM selection by summing token-level rewards up. PPO-S: PPO trained based on seq2seq RM, with sentence-level reward by summing token-level rewards up. PPO-T-Pos: PPO-T with only positive scores. RFT-S: RFT based on seq2seq RM selection by summing token-level rewards up. All methods are implemented based on the same code base, PKU-Alignment/SafeRLHF. For more details, including hyperparameter settings, please refer to our supplementary materials and appendix.
@@
Methods Length ($\Delta$downarrow) Win Rate $\uparrow$
@@
Table 1: Comparison of seq2seq RM with common alignment methods on Dialoguesum. $\Delta$ denotes the response length difference to Init-SFT. The Win Rate is the win-lose ratio compared to Init-SFT.
@@
The criterion for a preferred dialogue summary is its accuracy and conciseness in summarizing the dialogue content. However, our statistics reveal a bias: 61% of responses labeled as chosen in the synthetic Dialoguesum dataset are longer. This indicates a slight bias for length. As shown in Table 1, PPO is significantly affected by this long-response bias, generating obviously longer responses. In contrast, PPO-S and PPO-T, based on seq2seq RM, can effectively mitigate the long-response bias issue.
@@
Table 2: Comparison of RLHF based on seq2seq RM with common alignment methods on PKU-SafeRLHF. The Utility scores are derived from the RM from Dai et al. (2023), the MT Bench Score is the evaluation results from Zheng et al. (2024) and the Safety Win Rate is the win-lose ratio compared to Init-SFT, evaluated by GPT-4.
@@
The trade-off between safety and utility has been a long-term challenge in LLMs alignment (Bai et al. 2022). We empirically demonstrate that RLHF targeting safety preferences can easily induce LLMs to fall into a refusal-to-response paradigm, like “sorry but I cannot assist with that.”. We use the utility score model by Dai et al. (2023) and a comprehensive evaluation framework, MT Bench Zheng et al. (2024) to quantitatively present this metric. As illustrated in Table 2, PPO trained with the seq2scalar RM converges to a refusal-to-response paradigm with a low utility score, and it did not achieve the greatest safety improvement. This difference is particularly evident in the decrease in the utility and MT Bench scores and a reduction in variance, where variance indirectly reflects the diversity of generated content.
@@
Table 4: Comparison of Methods for Gemma-2B and Llama2-7B on OOD Xsum (Narayan, Cohen, and Lapata 2018) prompts, with RMs trained on TL;DR datasets.
@@
Models PKU-SafeRLHF Dialoguesum TL;DR
@@
Reward modeling is a crucial component of RLHF. Given that RM serves as imperfect proxies for human intentions, LLMs can exploit their vulnerabilities, exhibiting unexpected generalization to achieve high rewards during RL finetuning (Pitis 2023; Gao, Schulman, and Hilton 2023). It means RM fails to provide feedback that accurately aligns with human intentions: LLMs exploring unexpected generalization to obtain high rewards, which negatively impact their capabilities and fail to achieve alignment objectives (Tien et al. 2022; Knox et al. 2023; Lambert and Calandra 2023). Moskovitz et al. (2023) proposed the use of Lagrange multipliers to balance rewards representing various preferences, aiming to create a more nuanced reward structure. Similarly, Coste et al. (2023) explored ensembling multiple RMs to provide more accurate feedback by leveraging diverse perspectives. Ramé et al. (2024) involved averaging RM weights under different random settings to enhance robustness against variability. Additionally, Zhu, Jordan, and Jiao (2024) introduced iterative data refinement techniques to minimize noise within the reward signals. While these methods have demonstrated varying degrees of effectiveness, they all introduce additional models, training requirements, and hyperparameters, which may impact their robustness and scalability.
@@
Recent research on token level RLHF presents an impossible triangle of granularity, accuracy, and annotation cost. Chan et al. (2024) redistributed the seq2scalar RM’s output to each token based on attention values to achieve fine-grained feedback, but this still relies on the Bradley-Terry model’s accuracy over whole responses in learning preferences. Chen et al. (2024), from the perspective of minimum editing constraint, used a generative model to provide token-level reward feedback, but this requires an additional teacher model for correction annotations. Rafailov et al. (2024a) and Zeng et al. (2024) demonstrated the feasibility of achieving token-level alignment from a variant of RLHF, direct preference optimization (DPO) (Rafailov et al. 2024b). However, DPO’s offline and off-policy setting reduces its alignment efficiency and cannot be optimized when the preference dataset variability is insignificant (Tajwar et al. 2024). Xu et al. (2024) further showed through theoretical and experimental analysis that DPO is a suboptimal alternative to RLHF methods in various scenarios. This also raises concerns about using DPO as the reward model for RLHF by Zhong et al. (2024).

### v2

@@ Introduction
Large Language Models (LLMs) (achiam2023gpt; touvron2023llama) have significantly advanced capabilities in conversation, summarization, reasoning, and more tasks, leading to widespread application (meng2024application; anthropic2023claude). Alignment techniques (ji2023ai), particularly Reinforcement Learning from Human Feedback (RLHF) (ouyang2022training), have become crucial for aligning LLMs with human intentions and values, e.g., reducing toxicity (bai2022training; ji2024language; ji2024align). RLHF consists of two main steps: training a reward model (RM) based on binary human preferences and then fine-tuning the LLMs to maximize feedback from this RM (ouyang2022training; schulman2017proximal; languagefeedback). Current reward modeling implementations typically use the sequence-to-scalar (seq2scalar) approach, where a sequence, combined with a prompt and response, is scored as a scalar value (huang2024n+).
Figure 1: The traditional sequence-to-scalar reward model provides only coarse-grained scalar feedback on the LLM’s response, making it prone to exploiting unexpected generalization for high rewards during the RL fine-tuning phase (skalse2023misspecification), e.g., falling into a refusal-to-answer paradigm (bai2022training; dai2023safe). Without additional annotations, we propose a novel sequence-to-sequence reward modeling method that offers richer language feedback, improving RLHF performance.
@@
However, due to high divergence, noise, and sparsity in crowd-annotated preference datasets (wu2024fine), the traditional seq2scalar RM tends to capture patterns unrelated to genuine human preferences, leading to unexpected generalization (pan2021effects; gao2023scaling). It means RM fails to provide feedback that accurately aligns with human preference, leading RLHF to a deviation from actual alignment objective (di2022goal): LLMs obtain high rewards but negatively impact their capabilities and fail to achieve alignment objectives (tien2022causal; knox2023models; lambert2023alignment). For example, when the alignment objective is to reduce toxic responses, LLMs may learn the refusal-to-response paradigm instead of providing safe and helpful responses (dai2023safe; bai2022training). Existing methods require additional annotations to provide more comprehensive feedback, including regularization terms (moskovitz2023confronting; zhu2024iterative), models ensembling (coste2023reward), parameters combinations (rame2024warm) and additional model (dai2023safe), raising concerns about their scalability.
@@
  • Our further experiments demonstrate that the seq2seq RM not only mitigates unexpected behavior of RLHF but also enhances its alignment performance. Specifically, it achieves an average win rate of 76.9% in 3 NLP tasks, across models with 2B and 7B parameters.
@@
Given an LLM parameterized with $\theta$, we model its generation as a token-level Markov Decision Process (MDP), i.e., the tuple $\mathcal{T}=(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P})$. Here, let $t\in\mathcal{V}$ denote a token, where $\mathcal{V}$ represents the vocabulary set, and $L$ denotes the maximum number of tokens. The state space $\mathcal{S}$ can be defined as $\{(t_{1},t_{2},\ldots,t_{L})\mid t_{i}\in\mathcal{V}\}$  [^1] We define $t_{pad}$ as the padding token. When the length of a sentence is $l<L$, it will have LlL-l additional padding tokens.. The action space $\mathcal{A}=\mathcal{V}$ consists of all possible tokens in vocabulary. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns a reward value for generating a specific token given the current sequence of tokens. The transition probability function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the deterministic state transition model, concatenating the generated token to the existing sequence to form the next state. i.e., $s_{t+1}=(s_{t},a_{t})$.
@@
$\mathcal{L}_{\bm{\theta}}^{\text{RL}}=-\mathbb{E}_{a_{t}\sim\pi_{\bm{\theta}}(\cdot|s_{t})}[\sum_{t=0}^{T}(\mathcal{R}(s_{t},a_{t})+\beta\pi_{ref}(a_{t}|s_{t}))]$
where $\pi_{ref}$ is the reference distribution from the initial LLM, combined with coefficient $\beta$ to compute the KL divergence, preventing the LLM from deviating too much from the $\pi_{ref}$.
@@ Methods
The seq2scalar RM maps concatenated prompt and response pairs, denoted as $\bm{x},\bm{y^{l}}$ and $\bm{x},\bm{y^{w}}$, to distinct points in a scalar space. This makes it difficult for the token-level differences between responses to be distinguished in the scalar output. To enhance RM’s ability to discern fine-grained differences in responses, we propose augmenting this mapping from a scalar to a language space. Specifically, we extend the learning process from binary MLE in Equation 1 to SFT, a sequence MLE approach. This extension transforms the learning objective from scalar classification to sequence generation, achieving a higher distinction of responses, as shown in Equation Reward Modeling as Sequence MLE. Similar approaches have also been applied by aligner to make LLM aligners, which is an LLM designed to refine the output of preceding ones.
@@
Since the language feedback provided by the seq2seq reward model (RM) cannot be directly used for reinforcement learning (RL) optimization, we design a novel reward extraction mechanism that returns positive and negative scores as token-level feedback. Formally, let $\bm{x}$ denote the prompt used in RLHF, and $\bm{y}$ is the response generated by the actor LLM. We set $j$ as the first index where $r^{\text{seq}}(\bm{x},\bm{y})$ and $\bm{y}$ are not equal, with the subscript indicating the element at the corresponding position. The notation :j:j indicates the subsequence up to that index. Defining the inference function of the seq2seq RM as $r^{\text{seq}}(\cdot)$, our goal is to leverage $r^{\text{seq}}(\cdot)$ to obtain the token-level score $\mathcal{R}(\bm{x},\bm{y_{:i}})$ for all subsequence $\bm{y_{:i}}$ in $\bm{y}$. To achieve this, we first concatenate $\bm{x}$ and $\bm{y}$ and then input $\bm{y}$ token by token into the seq2seq RM. We positively reward tokens that are consistent with the seq2seq RM at the beginning and negatively reward those that become inconsistent thereafter. Figure 2 provides an intuitive example of this process.
@@
$\mathcal{R}(\bm{x},\bm{y_{:i-1}})=\begin{cases}1&\text{if }r^{\text{seq}}(\bm{x},\bm{y},\bm{y_{:i-1})}=\bm{y_{i}}\text{ and }i<j\\ -1&\text{until divergence}.\end{cases}$
@@
We conducted experiments on the popular open-source LLMs Gemma-2B (team2024gemma) and Llama2-7B (touvron2023llama). The selected tasks included single-turn safe dialogue, dialogue summarization, and text summarization. For single-turn safe dialogue, we chose the PKU-SafeRLHF dataset (ji2024beavertails; ji2024pku). To verify the performance of our method on human preference and synthetic preference settings, we used the Dialogsum (chen2021dialogsum) for the dialogue summarization task and the TL;DR (stiennon2020learning; volske-etal-2017-tl) for text summarization. For the former, we employed GPT-4 for preference annotation, while the latter already contains human-annotated preferences.
@@
Our evaluation primarily uses GPT-4 evaluation and phenomenon analysis for specific tasks. In single-turn safety dialogue tasks, we analyze whether LLMs fall into an overly conservative refusal-to-response paradigm, leading to a significant drop in utility. Therefore, we use the helpfulness reward model scores provided by dai2023safe and MT Bench score (zheng2024judging) as the metric of utility. In text summarization and dialogue summarization tasks, we refer to (stiennon2020learning) and use accuracy, conciseness, and information consistency as evaluation criteria. To decrease randomness, we ran the above evaluation over 3 times.
@@
Our experiment includes comparisons between popular baselines and our proposed variants. We define their abbreviations as follows. Baselines. Init-SFT: Pre-trained model finetuned for the corresponding task, also used as the initial model for other alignment methods. PPO: Vanilla PPO trained based on seq2scalar RM (ouyang2022training; schulman2017proximal). DPO: Direct Preference Optimization (rafailov2024direct). RFT: Rejection sampling fine-tuning (touvron2023llama) based on seq2scalar RM selection over 8 responses generated by the Init-SFT. Variants. PPO-T: PPO trained based on seq2seq RM, with token-level reward. RFT-S: RFT based on seq2seq RM selection by summing token-level rewards up. PPO-S: PPO trained based on seq2seq RM, with sentence-level reward by summing token-level rewards up. PPO-T-Pos: PPO-T with only positive scores. All methods are implemented based on the same code base, PKU-Alignment/SafeRLHF. For more details, including hyperparameter settings, please refer to our supplementary materials and appendix.
@@
Methods Length ($\Delta$ $\downarrow$) Win Rate $\uparrow$
@@
Table 1: Comparison of seq2seq RM with common alignment methods on Dialogsum. $\Delta$ denotes the response length difference to Init-SFT. The Win Rate is the win-lose ratio compared to Init-SFT judged by GPT-4.
@@
The criterion for a preferred dialogue summary is its accuracy and conciseness in summarizing the dialogue content. However, our statistics reveal a bias: 61% of responses labeled as chosen in the synthetic Dialogsum dataset are longer. This indicates a slight bias for length. As shown in Table 1, PPO is significantly affected by this long-response bias, generating obviously longer responses. In contrast, PPO-S and PPO-T, based on seq2seq RM, can effectively mitigate the long-response bias issue.
@@
Table 2: Comparison of RLHF based on seq2seq RM with common alignment methods on PKU-SafeRLHF. The Utility scores are derived from the RM from dai2023safe, the MT Bench Score is the evaluation results from zheng2024judging and the Safety Win Rate is the win-lose ratio compared to Init-SFT, evaluated by GPT-4.
@@
The trade-off between safety and utility has been a long-term challenge in LLMs alignment (bai2022training). We empirically demonstrate that RLHF targeting safety preferences can easily induce LLMs to fall into a refusal-to-response paradigm, like “sorry but I cannot assist with that.”. We use the utility score model by dai2023safe and a comprehensive evaluation framework, MT Bench zheng2024judging to quantitatively present this metric. As illustrated in Table 2, PPO trained with the seq2scalar RM converges to a refusal-to-response paradigm with a low utility score, and it did not achieve the greatest safety improvement. This difference is particularly evident in the decrease in the utility and MT Bench scores and a reduction in variance, where variance indirectly reflects the diversity of generated content.
@@
Table 4: Comparison of methods for Gemma-2B and Llama2-7B on OOD Xsum (narayan2018don) prompts, with RMs trained on TL;DR datasets.
@@
Models PKU-SafeRLHF Dialogsum TL;DR
@@
Reward modeling is a crucial component of RLHF. Given that RM serves as imperfect proxies for human intentions, LLMs can exploit their vulnerabilities, exhibiting unexpected generalization to achieve high rewards during RL finetuning (pitis2023failure; gao2023scaling). It means RM fails to provide feedback that accurately aligns with human intentions: LLMs exploring unexpected generalization to obtain high rewards, which negatively impact their capabilities and fail to achieve alignment objectives (tien2022causal; knox2023models; lambert2023alignment). moskovitz2023confronting proposed the use of Lagrange multipliers to balance rewards representing various preferences, aiming to create a more nuanced reward structure. Similarly, coste2023reward explored ensembling multiple RMs to provide more accurate feedback by leveraging diverse perspectives. rame2024warm involved averaging RM weights under different random settings to enhance robustness against variability. Additionally, zhu2024iterative introduced iterative data refinement techniques to minimize noise within the reward signals. While these methods have demonstrated varying degrees of effectiveness, they all introduce additional models, training requirements, and hyperparameters, which may impact their robustness and scalability.
@@
Recent research on token level RLHF presents an impossible triangle of granularity, accuracy, and annotation cost. chan2024dense redistributed the seq2scalar RM’s output to each token based on attention values to achieve fine-grained feedback, but this still relies on the Bradley-Terry model’s accuracy over whole responses in learning preferences. chen2024improving, from the perspective of minimum editing constraint, used a generative model to provide token-level reward feedback, but this requires an additional teacher model for correction annotations. rafailov2024r and zeng2024token demonstrated the feasibility of achieving token-level alignment from a variant of RLHF, direct preference optimization (DPO) (rafailov2024direct). However, DPO’s offline and off-policy setting reduces its alignment efficiency and cannot be optimized when the preference dataset variability is insignificant (tajwar2024preference). xu2024dpo further showed through theoretical and experimental analysis that DPO is a suboptimal alternative to RLHF methods in various scenarios. This also raises concerns about using DPO as the reward model for RLHF by zhong2024dpo.